# How to Tell When a Result Will Replicate: Significance and Replication in Distributional Null Hypothesis Tests

### v2

@@ 4.4 Comparing predicted and observed replication probabilities
To assess the distributional model of replication, for each of the $16$ tasks we generated every possible pair of experiments from two different US sites and calculated the replication probability $p_{rep}$ for the second or target experiment in each pair based on the result $t$ in the first or predictor experiment. This calculation used the between-experiment variance of means for the task overall, $S_{0}^{2}$, the within-experiment variance in the predictor experiment $S^{2}$ and the sample size of the predictor experiment $N$, and, given these values, generating the probability of replication for an experiment with the sample size $N_{r}$ of the paired target experiment. This calculation assumed $c=1$ and so did not make use of the sample variance of the target experiment (and so represents the calculation a researcher would carry out when they have observed a surprising result in an experiment with sample size $N$ and within-experiment variance $S^{2}$, and are estimating the probability of replicating that result in a planned replication with sample size $N_{r}$ but unknown within-experiment variance). We calculated this replication probability for a range of commonly used significance levels $\alpha=,0.05,0.025,0.01,0.005,0.001$. For each target experiment and each level $\alpha$ we compared this predicted replication probability against a dichotomous successful replication variable ($p_{sig}\leq\alpha$) which indicated whether the target experiment did, in fact, replicate successfully at level $\alpha$.
As an an initial check we calculated, for each pair of experiments in a given task, both the closed-form approximation $p_{rep}$ and the double-integral expression $p_{rep}(t|\hat{b})$. As before, these were almost identical (correlation $r=0.94$) supporting the closed-form approximations, and so we primarily make use of the closed-form approximation $p_{sig}$ in our analysis (Figure $3$ gives a visual comparison of $p_{rep}$ and $p_{rep}(t|\hat{b})$ values).
@@ 5 Comparing with Bayesian approaches
This distributional model mixes ‘frequentist-style’ hypothesis testing with ‘Bayesian-style’ updating and posterior prediction for estimating replication probabilities. Advocates of the Bayesian approach may consider that, in the words of one reviewer, we are ‘teetering on the edge of Bayes already’ and ask why we do not ‘go all the way’, by abandoning tests of statistical significance and moving to alternative Bayesian approaches. In this section we address this possibility by comparing the distributional model against the best-known and most commonly used Bayesian alternative to significance testing, the Bayesian or JZS $t$-test (Rouder et al.,, 2009).
5.1 JZS Bayesian $t$-test
@@
To use a Bayes Factor approach to assess evidence for or against a null hypothesis, the JZS Bayesian $t$-test (Rouder et al.,, 2009) takes the parameter of interest to be the sample mean $\overline{X}$ and takes $H_{0}$ to be the standard point-form null
@@
The JZS Bayesian $t$-test is implemented in the R BayesFactor package via the function ‘ttest.tstat’ (Morey and Rouder,, 2021). To compare this approach against our distributional null proposal, we used this package to calculate $BF_{10}$ for every experiment in the Many-Labs dataset, and compared these $BF_{10}$ values against the distributional significance $p_{sig}$ for that result on a $\log_{10}$ scale (Figure 2(Right)). This comparison gave results almost identical those seen in the earlier point-form $p$ comparison. There were many extremely high $BF_{10}$ values, with $33\%$ of experiments having $BF_{10}>10^{5}$ and $17\%$ of experiments having $BF_{10}>10^{15}$. Distributional $p_{sig}$ values were again orders of magnitude more conservative. Many results gave strong or very strong evidence against the null in the Bayesian test while not reaching the minimum level of significance in the distributional test: $20\%$ of results with $BF_{10}>30$ (very strong evidence against the null) had $p_{sig}>0.05$. Even results with apparently very strong evidence against the null (in the Bayesian $t$-test) turn out to be consistent with the null hypothesis when between-experiment variation is taken into account.
@@
The left and right scatterplots in Figure 2 are extremely similar, suggesting that the logs of the point-form $p$ and the Bayes Factor $BF_{10}$ are related; or more precisely (since Figure 2(Left) has a negative base-10 log scale on the $y$ axis) that $\log_{10}(BF_{10})$ and $\log_{10}(1/p)$ are related. A comparison of values of $\log_{10}(BF_{10})$ versus $\log_{10}(1/p)$ across experiments and tasks in the Many Labs dataset and confirms that these measures have an almost perfect linear relationship, with a correlation of $r=0.9999$ and with the line of best fit being $\log_{10}(BF_{10})=0.99\log_{10}(1/p)-1.59$. This means that the point-form and the Bayesian $t$-test convey essentially the same information, at least in this dataset. This close relationship between the point-form $p$ value and the Bayesian $t$-test measure is not surprising: it has long been observed empirically that Bayes Factor measures involving the point-form null and the point-form $p$-value measure are strongly correlated (e.g. Wetzels et al.,, 2011) and that decisions based on these two measures are often equivalent (to quote Jeffreys: “As a matter of fact I have applied my significance tests to numerous applications that have also been worked out by Fisher’s, and have not yet found a disagreement in the actual decisions reached”; cited in Ly et al.,, 2016). Given this relationship it seems clear that the various problems with point-form hypothesis testing (failures of replication for significant results, the fact that the point-form null hypotheses are always false, the effects of sample size on significance, and so on) will also apply to Bayesian approaches that involve the point-form null: we argue that these problems can be addressed, not by moving from a Frequentist to a Bayesian position, but instead by moving from point-form to distributional nulls.
@@ 6 Conclusions
A third possible objection concerns the availability of alternative hypothesis-testing methods, such as the Bayesian $t$-test (commonly put forward as a replacement for the significance testing approach). ‘While significance testing was appropriate in the last century’, we imagine the objection goes, ‘today we have better statistical methods; we don’t need to test for significance’. Our response here is to note that measures of significance in the point-form $t$-test and measures of evidence in the JZS Bayesian $t$-test appear to be almost exactly equivalent in order of magnitude, at least in the Many Labs $1$ data, and so replacing significance testing against a point-form null with this form of Bayesian $t$-test may not, therefore, represent a meaningful advance.
While useful, our distributional approach is subject to various caveats. First, our approach assumes normally distributed variation or error both within and between experiments. This assumption is reasonable for simple experiments involving one or two variables ($t$-tests, tests of correlation or contingency) but does not hold in more complex experimental designs involving multiple variables, such as ANOVAs, multiple regression analyses, and mixed models. This is because the internal structure in such designs can cause random between-experiment differences to have systematic non-normal effects on between-experiment variability. Applying our distributional approach to between-experiment variation in this more complex setting would be challenging, requiring a statistical model of between-experiment variance in each component of the experimental design and how that variance affects the overall between-experiment distribution of results.

### v3

@@ 4.4 Comparing predicted and observed replication probabilities
To assess the distributional model of replication, for each of the $16$ tasks we generated every possible pair of experiments from two different US sites and calculated the replication probability $p_{rep}$ for the second or target experiment in each pair based on the result $t$ in the first or predictor experiment. This calculation used the between-experiment variance of means for the task overall, $S_{0}^{2}$, the within-experiment variance in the predictor experiment $S^{2}$ and the sample size of the predictor experiment $N$, and, given these values, generating the probability of replication for an experiment with the sample size $N_{r}$ of the paired target experiment. This calculation assumed $c=1$ and so did not make use of the sample variance of the target experiment (and so represents the calculation a researcher would carry out when they have observed a surprising result in an experiment with sample size $N$ and within-experiment variance $S^{2}$, and are estimating the probability of replicating that result in a planned replication with sample size $N_{r}$ but unknown within-experiment variance). We calculated this replication probability for a range of commonly used significance levels $\alpha=0.05,0.025,0.01,0.005,0.001$. For each target experiment and each level $\alpha$ we compared this predicted replication probability against a dichotomous successful replication variable ($p_{sig}\leq\alpha$) which indicated whether the target experiment did, in fact, replicate successfully at level $\alpha$.
As an initial check we calculated, for each pair of experiments in a given task, both the closed-form approximation $p_{rep}$ and the double-integral expression $p_{rep}(t|\hat{b})$. As before, these were almost identical (correlation $r=0.94$) supporting the closed-form approximations, and so we primarily make use of the closed-form approximation $p_{sig}$ in our analysis (Figure $3$ gives a visual comparison of $p_{rep}$ and $p_{rep}(t|\hat{b})$ values).
@@ 5 Comparing with Bayesian approaches
This distributional model mixes ‘frequentist-style’ hypothesis testing with ‘Bayesian-style’ updating and posterior prediction for estimating replication probabilities. Advocates of the Bayesian approach may consider that we are ‘teetering on the edge of Bayes already’ and ask why we do not go ‘all the way’, by abandoning tests of statistical significance and moving to alternative Bayesian approaches. In this section we address this possibility by comparing the distributional model against the best-known and most commonly used Bayesian alternative to significance testing, the Bayesian or JZS $t$ test (Rouder et al.,, 2009).
5.1 JZS Bayesian $t$ test
@@
To use a Bayes Factor approach to assess evidence for or against a null hypothesis, the JZS Bayesian $t$ test (Rouder et al.,, 2009) takes the parameter of interest to be the sample mean $\overline{X}$ and takes $H_{0}$ to be the standard point-form null
@@
The JZS Bayesian $t$ test is implemented in the R BayesFactor package via the function ‘ttest.tstat’ (Morey and Rouder,, 2021). To compare this approach against our distributional null proposal, we used this package to calculate $BF_{10}$ for every experiment in the Many-Labs dataset, and compared these $BF_{10}$ values against the distributional significance $p_{sig}$ for that result on a $\log_{10}$ scale (Figure 2(Right)). This comparison gave results almost identical to those seen in the earlier point-form $p$ comparison. There were many extremely high $BF_{10}$ values, with $33\%$ of experiments having $BF_{10}>10^{5}$ and $17\%$ of experiments having $BF_{10}>10^{15}$. Distributional $p_{sig}$ values were again orders of magnitude more conservative. Many results gave strong or very strong evidence against the null in the Bayesian test while not reaching the minimum level of significance in the distributional test: $20\%$ of results with $BF_{10}>30$ (very strong evidence against the null) had $p_{sig}>0.05$. Even results with apparently very strong evidence against the null (in the Bayesian $t$ test) turn out to be consistent with the null hypothesis when between-experiment variation is taken into account.
@@
The left and right scatterplots in Figure 2 are extremely similar, suggesting that the logs of the point-form $p$ and the Bayes Factor $BF_{10}$ are related; or more precisely (since Figure 2(Left) has a negative base-10 log scale on the $y$ axis) that $\log_{10}(BF_{10})$ and $\log_{10}(1/p)$ are related. A comparison of values of $\log_{10}(BF_{10})$ versus $\log_{10}(1/p)$ across experiments and tasks in the Many Labs dataset and confirms that these measures have an almost perfect linear relationship, with a correlation of $r=0.9999$ and with the line of best fit being $\log_{10}(BF_{10})=0.99\log_{10}(1/p)-1.59$. This means that the point-form and the Bayesian $t$ test convey essentially the same information, at least in this dataset. This close relationship between the point-form $p$ value and the Bayesian $t$ test measure is not surprising: it has long been observed empirically that Bayes Factor measures involving the point-form null and the point-form $p$-value measure are strongly correlated (e.g. Wetzels et al.,, 2011) and that decisions based on these two measures are often equivalent (to quote Jeffreys: “As a matter of fact I have applied my significance tests to numerous applications that have also been worked out by Fisher’s, and have not yet found a disagreement in the actual decisions reached”; cited in Ly et al.,, 2016). Given this relationship it seems clear that the various problems with point-form hypothesis testing (failures of replication for significant results, the fact that the point-form null hypotheses are always false, the effects of sample size on significance, and so on) will also apply to Bayesian approaches that involve the point-form null: we argue that these problems can be addressed, not by moving from a Frequentist to a Bayesian position, but instead by moving from point-form to distributional nulls.
@@ 6 Conclusions
A third possible objection concerns the availability of alternative hypothesis-testing methods, such as the Bayesian $t$ test (commonly put forward as a replacement for the significance testing approach). ‘While significance testing was appropriate in the last century’, we imagine the objection goes, ‘today we have better statistical methods; we don’t need to test for significance’. Our response here is to note that measures of significance in the point-form $t$ test and measures of evidence in the JZS Bayesian $t$ test appear to be almost exactly equivalent in order of magnitude, at least in the Many Labs $1$ data, and so replacing significance testing against a point-form null with this form of Bayesian $t$ test may not, therefore, represent a meaningful advance.
While useful, our distributional approach is subject to various caveats. First, our approach assumes normally distributed variation or error both within and between experiments. This assumption is reasonable for simple experiments involving one or two variables ($t$ tests, tests of correlation or contingency) but does not hold in more complex experimental designs involving multiple variables, such as ANOVAs, multiple regression analyses, and mixed models. This is because the internal structure in such designs can cause random between-experiment differences to have systematic non-normal effects on between-experiment variability. Applying our distributional approach to between-experiment variation in this more complex setting would be challenging, requiring a statistical model of between-experiment variance in each component of the experimental design and how that variance affects the overall between-experiment distribution of results.